OpenStudy (anonymous):

3.Substitute your birthday into the equation Square root of the quantity x minus y + m = d, where y is the last two digits of your birth year, m is the month, and d is the day. If you were born on 7/10/1856 like Nikola Tesla, your equation would be Square root of the quantity x minus 56 + 7 = 10. Solve for x and identify if it is an extraneous solution. My birthday is 06/06/1998 who ever answers this I will give a medal to.

OpenStudy (chrisdbest):

Dang my birthday is May 9th 1997! This seems a little tricky

OpenStudy (anonymous):

I know I tried asking my teacher but he wouldn't help me .-.

OpenStudy (astrophysics):

What's stopping you from doing what the example showed you?

OpenStudy (chrisdbest):

So I got \[\sqrt{x-98} +6=6\]

OpenStudy (chrisdbest):

But I don't know if the question is asking for this: \[\sqrt{x-98} +6=6\] or this: \[\sqrt{x-98+6}=6\]

OpenStudy (anonymous):

I got sqrt x-98 +6=6 I just don't understand how to solve the equation

OpenStudy (astrophysics):

Ok, but answer what chris has asked you, please specify your equation.

OpenStudy (astrophysics):

Is it \[\sqrt{x-98}+6=6 ~~~~ \text{or}~~~\sqrt{x-98+6}=6\]...

OpenStudy (anonymous):

\[\sqrt{x-98} +6=6\] is what I got

OpenStudy (astrophysics):

Ok good, so now you want to look for the extraneous solution

OpenStudy (astrophysics):

So we want to solve for x right?

OpenStudy (anonymous):

Yes wouldn't we square the x-98 to get rid of the square root?

OpenStudy (astrophysics):

Right! See you know this stuff :)

OpenStudy (chrisdbest):

And what's next?

OpenStudy (astrophysics):

First lets simplify it a bit, move the 6 on the right

OpenStudy (astrophysics):

Then we'll take the square root

OpenStudy (anonymous):

So we take the square root of the 6 right?

OpenStudy (chrisdbest):

Yes, so your left with x-98 + 6 = 6

OpenStudy (astrophysics):

Well I meant, lets make it easier on ourselves by doing this \[\sqrt{x-98}-6=6 \implies \sqrt{x-98} = 12\]

OpenStudy (chrisdbest):

You know what to do from here?

OpenStudy (anonymous):

Yes thank you guys so much for the help!

OpenStudy (astrophysics):

Np, everything make sense now?